Circular No. 8388

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dv IN MCG –01-6-12 Further to IAUC 8387, M. Ganeshalingam and W. Li report the LOSS

discovery of an apparent supernova (mag ∼ 18.5) on an unfiltered KAIT image taken on Aug. 13.49 UT. The new object was also marginally detected on earlier images taken on July 21.49, 27.45, and Aug. 8.44 (all with limiting mag ∼ 18.5–19.0). SN 2004dv is located at α = 1^{h}58^{m}52^{s}.76, δ = −8^{o}10′11′′.5 (equinox 2000.0), which is 12′′.5 east and 17′′.5 south of the nucleus of MCG –01-6-12. A KAIT image taken on 2003 Sept. 11.36 showed nothing at this position (limiting mag ∼ 20.0).

SUPERNOVAE 2004dj AND 2004du W. Li and A. V. Filippenko, University of California, Berkeley (UCB);

and S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology, report: “We have identified a progenitor candidate for the type II-P supernova 2004dj (cf. IAUC 8377, 8378, 8384, 8385) in multiband images obtained at the Nordic Optical Telescope (NOT) by Larsen and Richtler (1999, A.Ap. 345, 59), under good (0′′.8) seeing conditions. This star, with V = 21.1 and V − I = +1.4, is 0′′.3 southwest of the radio position of the supernova (cf. IAUC 8379). Assuming little reddening toward the supernova, other than Galactic (Schlegel et al. 1998, Ap.J. 500, 525), and assuming a true distance modulus of 27.48 mag (Freedman et al. 2001, Ap.J. 553, 47), this corresponds to an extinction-corrected absolute V magnitude of −6.5 and an unreddened V − I color of +1.3. The color and magnitude are consistent with a red supergiant with initial mass of ∼ 17 M_{⊙}. A brighter object to the northeast, Sandage’s ‘star’ 96 (cf. IAUC 8385), is within 0′′.5 of the reported optical position of SN 2004dj (cf. IAUC 8377) but is 1′′.0 west and 1′′.2 south of the radio position. The profile of ‘star’ 96 appears nonstellar in the I-band NOT image. We consider it less likely that ‘star’ 96 is the progenitor of this supernova and suggest that definitive identification of the progenitor will only be possible when high-resolution images of SN 2004dj become available.”

Filippenko, R. J. Foley (UCB), and Li add that inspection of CCD spectra (range 320–930 nm), obtained by C. Fassnacht and J. McKean (University of California, Davis) on Aug. 13 UT with the Keck-I 10-m telescope (+ LRIS), shows that SN 2004du (IAUC 8387) is a young type-IIn supernova. Very weak, narrow Hα emission is superposed on a featureless, extremely blue continuum. There is a hint of Hβ and He I 587.6-nm emission, as well.

2004 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green